### v1

# Understanding Practical Membership Privacy of Deep Learning

Abstract
We apply a state-of-the-art membership inference attack (MIA) to systematically test the practical privacy vulnerability of fine-tuning large image classification models. We focus on understanding the properties of data sets and samples that make them vulnerable to membership inference. In terms of data set properties, we find a strong power law dependence between the number of examples per class in the data and the MIA vulnerability, as measured by true positive rate of the attack at a low false positive rate. For an individual sample, large gradients at the end of training are strongly correlated with MIA vulnerability.
1 Introduction
Machine learning models are prone to memorising their training data, which makes them vulnerable to privacy attacks such as membership inference attacks (MIAs; Shokri et al., 2017; Carlini et al., 2022) and reconstruction attacks (e.g. Balle et al., 2022; Nasr et al., 2023). Differential privacy (DP; Dwork et al., 2006b) provides protection against these attacks, but strong formal protection often comes at the cost of significant loss of model utility.
Finding the correct balance between making models resistant to attacks while maintaining a high utility is important for many applications. In health, for example, many European countries and soon also the EU within the European Health Data Space have requirements that models trained on health data that are made publicly available must be anonymous, i.e. they must not contain information that can be linked to an identifiable individual. On the other hand, loss of utility of the model may compromise the health benefits that might be gained from it.
In this paper, our aim is to apply a state-of-the-art MIA systematically to help understand practical privacy risks training deep-learning-based classifiers without DP protections. Our case study focuses on understanding and quantifying factors that influence the vulnerability of non-DP deep learning models to MIA. Since the earliest MIA methods, there has been evidence that classification models with more classes are more vulnerable to MIA (Shokri et al., 2017). Similarly, there is evidence that models trained on fewer samples can be more vulnerable (Chen et al., 2020; Németh et al., 2023). Based on extensive experimental data over many data sets with varying sizes, we uncover a power law describing the vulnerability remarkably well.
Previously Tobaben et al. (2023) and Yu et al. (2023) have presented limited studies of how data set properties affect the MIA vulnerability. Tobaben et al. (2023) reported how the MIA vulnerability of few-shot image classification is affected by the number of shots (i.e. the number of examples available per class). Yu et al. (2023) studied how the MIA vulnerability of a class in an image classifier trained from scratch depends on the average individual privacy parameters of that class. Our work significantly expands on both of these by more systematic study and modelling of data set properties and by providing a more detailed look at individual samples, not just classes.
List of contributions
In this work we perform an extensive empirical study on the MIA vulnerability of deep learning models. We will focus on transfer learning setting on computer vision tasks, where a large pre-trained neural network is fine-tuned on a sensitive data set, and the MIA is performed on the final classifier. We find that the MIA vulnerability has a strong correlation with certain data set properties (Section 5) and make the following contributions:
  1. 1.
    Power law: We uncover a power law between the number of examples per class and the vulnerability to MIA (tpr at fixed low fpr) based on extensive experimental data over many data sets with varying sizes (See Figure 1(a)).
  2. 2.
    Regression model: We utilize the observations to train a regression model to predict MIA vulnerability (tpr at fixed low fpr) based on examples per class ($S$) and number of classes ($C$) and show both very good fit on the training data as well as good prediction quality on unseen data from a different feature extractor (See Figure 3).
Additionally, in Section 6 we extend the analysis to individual samples and discover a correlation between vulnerability of individual samples and the gradient norms during training. We further study the correlation between sample’s vulnerability and the similarity to the other samples in the feature space.
2 Background
Notation for the properties of the training data set $\mathcal{\mathcal{D}}$:
  • $C$ for the number of classes
  • $S$ for shots (examples per class)
  • $\mathcal{|\mathcal{D}|}$ for training data set size ($\mathcal{|\mathcal{D}|}=CS$)
We denote the number of MIA shadow models with $M$.
Membership inference attacks (MIAs) aim to infer whether a particular sample was part of the training set of the targeted model Shokri et al. (2017). Thus, they can be used to determine lower bounds on the privacy leakage of models to complement the theoretical upper bounds obtained through differential privacy.
Likelihood Ratio attack (LiRA; Carlini et al., 2022) While many different MIAs have been proposed Hu et al. (2022), in this work we consider the SOTA: the Likelihood Ratio Attack (LiRA). LiRA assumes an attacker that has black-box access to the attacked model, knows the training data distribution, the training set size, the model architecture, hyperparameters and training algorithm. Based on this information, the attacker can train so-called shadow models (Shokri et al., 2017) which imitate the model under attack but for which the attacker knows the training data set.
LiRA exploits the observation that the value of the loss function used to train a model is often lower for examples that have been part of the training set than for examples that have not been. For a target sample $x$, LiRA trains the shadow models: 1 with $x$ as a part of the training set ($x\in\mathcal{D}$) 2 and without $x$ in the training set ($x\notin\mathcal{D}$) . After training the shadow models, $x$ is passed through the shadow models, and based on the losses (or predictions) two Gaussian distributions are formed, one for the losses of $x\in\mathcal{D}$ shadow models and one for the $x\notin\mathcal{D}$. Finally the attacker computes the loss for the point $x$ using the model under attack and determines using a likelihood ratio test on the distributions built from the shadow models whether it is more likely that $x\in\mathcal{D}$ or $x\notin\mathcal{D}$. Carlini et al. (2022) proposed an optimization for performing LiRA for multiple models and points without training a computationally infeasible amount of shadow models. We utilize this optimization in our experiments.
In Appendix A we provide background on Differential privacy (DP; Dwork et al., 2006b) and DP stochastic gradient descent (DP-SGD) that we use for training DP models.
3 Methods
In this Section, we introduce the preliminaries for evaluating MIA vulnerability.
3.1 Measuring MIA vulnerability
Using the likelihood-ratios from LiRA Carlini et al. (2022) as a score, we can build a binary classifier to predict whether sample belongs to the training data or not. The operational characteristics of such classifier can be used to measure the success of the MIA. More specifically, throughout the rest of the paper, we will use the true positive rate (tpr) at a specific false positive rate (fpr) as a measure for the vulnerability. Identifying even a small number of samples with high confidence is considered harmful (Carlini et al., 2022) and thus we focus on the regions of small fpr.
The tpr and fpr can be also connected to DP. Kairouz et al. (2015) have shown in Equation 1 that any classifier distinguishing the training samples based on the results of a DP algorithm has an upper bound for the tpr that depends on the privacy parameters $(\epsilon,\delta)$. Since MIA is trying to build exactly these types of classifiers, we can use the upper bounds by Kairouz et al. (2015) to validate the privacy claims, and also to better understand the gap between the theoretical privacy guarantees and the realistic attacks.
Theorem 1 (Kairouz et al. (2015)).
A mechanism $\mathcal{M}:\mathcal{X}\to\mathcal{Y}$ is $(\epsilon,\delta)$-DP if and only if for all adjacent $\mathcal{D}\sim\mathcal{D}^{\prime}$
$\mathrm{TPR}\leq\min\{e^{\epsilon}\mathrm{FPR}+\delta,1-e^{-\epsilon}(1-\delta-\mathrm{FPR})\}\ .$ (1)
3.2 Measuring the uncertainty for TPR
The tpr values from the LiRA based classifier can be seen as maximum likelihood-estimators for the probability of producing true positives among the positive samples. Since we have a finite number of samples for our estimation, it is important to estimate the uncertainty in these estimator. Therefore, when we report the tpr values for single repeat of the learning algorithm, we estimate the stochasticity of the tpr estimate by using Clopper-Pearson intervals Clopper & Pearson (1934). Given tp true positives among p positives, the $1-\alpha$ confidence Clopper-Pearson interval for the tpr is given as
$\begin{split}&B(\alpha/2;\textsc{tp},\textsc{p}-\textsc{tp}+1)<\textsc{tpr}\\ &\textsc{tpr}<B(1-\alpha/2;\textsc{tp}+1,\textsc{p}-\textsc{tp}),\end{split}$ (2)
where $B(q;a,b)$ is the $q$th-quantile of $\text{Beta}(a,b)$ distribution.
4 Experiment overview
Throughout the experiments, we utilize LiRA to understand the vulnerability of the trained models.
We fine-tune pre-trained models on sensitive downstream data sets and assess the vulnerability using LiRA. We use the following setup for all experiments:
  • Pre-trained models. BiT-M-R50x1 (R-50; Kolesnikov et al., 2020) with 23.5M parameters and Vision Transformer ViT-Base-16 (ViT-B; Dosovitskiy et al., 2021) with 85.8M parameters, both pre-trained on the ImageNet-21K data set (Russakovsky et al., 2015)
  • Parameterization. Due to computational cost we only consider fine-tuning subsets of the model parameters. We consider the following configurations: 1 Head: training a linear last layer classifier on top of the pre-trained model 2 FiLM: Head and additionally training parameter-efficient FiLM (Perez et al., 2018) adapters scattered throughout the network (see Section B.1)
  • Fine-tuning data sets. We utilize different fine-tuning data sets (see Section B.3 for details) with different $S$, $C$ and $|\mathcal{D}|$ as specified in the experiments.
  • Hyperparameter optimization (HPO). We utilize the library Optuna Akiba et al. (2019) with the Tree-structured Parzen Estimator (TPE; Bergstra et al., 2011) sampler with 20 iterations. We optimize the batch size, learning rate and number of epochs. Additionally for DP-SGD we optimize the clipping norm. Details in Section B.2.
  • DP-SGD. We are using Opacus Yousefpour et al. (2021) and compute the noise multiplier based on the target ($\epsilon,\delta$).
  • DP Accounting. We use numerical privacy accounting for accurate evaluation of the cumulative privacy loss of the optimisation (Koskela et al., 2020, 2021) using the PRV accountant (Gopi et al., 2021).
  • LiRA. We base our experiments on the publically available source code of Carlini et al. (2022). We utilize $M=256$ shadow models unless specified otherwise.
We illustrate the effectiveness of LiRA and the effect of different levels of DP on the MIA vulnerability (tpr at $\textsc{fpr}=0.001$) in Figure 1. It shows the MIA vulnerability as a function of privacy budget (additional versions for different fprs in Figure A.1) and two upper bounds on the tpr at fpr emerging from the theoretical guarantees of DP-SGD.
Figure 1: MIA vulnerability (tpr at $\textsc{fpr}=0.001$) as a function of privacy budget ($\epsilon$ at $\delta=10^{-5}$) when attacking different backbones fine-tuned on CIFAR10 and CIFAR100 at $S\in\{25,50\}$. The solid line displays the median and the error bars the minimum of the lower bounds and the maximum of the upper bounds for the Clopper-Pearson CIs over 15 seeds. The upper bound is the theoretical DP bound based on Kairouz et al. (2015). We present the bound computed with $\delta=10^{-5}$ as well as the optimal bound based on multiple $\delta$.
We compute the upper bounds using the method of Kairouz et al. (2015) in Equation 1. Simply using $\delta=10^{-5}$ and the associated $\epsilon$ gives a relatively weak bound, that can be sharpened considerably by finding an optimal $\delta$ and corresponding $\epsilon(\delta)$, which is made possible by the particular mechanism deployed, that gives the tightest possible bound. For the attacked few-shot models the bounds are quite tight for the medium privacy budget of $\epsilon=1$ but very loose at low privacy budgets of $\epsilon=8$ and thus meaningless for low privacy levels.
5 Predicting MIA vulnerability for data sets
In this Section, we investigate how different properties of data sets affect the MIA vulnerability in the non-DP setting. Based on our observations, we propose a method to predict the vulnerability to MIA using these properties of the data set.
We base our experiments on a subset of the few-shot benchmark VTAB  (Zhai et al., 2019) that achieves a classification accuracy $>80\%$ and thus would considered to be used by practitioners. For more detailed descriptions of the data sets, see Table A2. In this Section, we will focus on a fine-tuned last layer classifier (Head) trained on top of a ViT-B, pre-trained on ImageNet-21k Russakovsky et al. (2015). The results for using R-50 as a backbone can be found in Section C.1.
5.1 Correlations with data set properties
Using the setting described above, we study how the number of classes and the number of shots affect the vulnerability (tpr at fpr as described in Section 3.1). We make the following observations:
  • A larger number of $S$ (shots) decrease the vulnerability in a power law relation as demonstrated in Figure 1(a). We provide further evidence of this in Figures A.2, A3 and A4 in the Appendix.
  • Contrary, a larger number of $C$ (classes) increases the vulnerability as demonstrated in Figure 1(b) with further evidence in Figures A.3, A5 and A6 in the Appendix. However, the trend w.r.t. $C$ is not as clear as with $S$.
(a) as function of $S$ (shots)
(b) as function of $C$ (classes)
Figure 2: MIA vulnerability (tpr at $\textsc{fpr}=0.001$) as a function of dataset properities when attacking a ViT-B Head fine-tuned without DP on different data sets. The solid line displays the median and the error bars the minimum of the lower bounds and maximum of the upper bounds for the Clopper-Pearson CIs over multiple seeds (6 for Figure 1(a) and 12 for Figure 1(b))
5.2 Model to predict data set vulnerability
Based on the trends seen in Figures 1(b) and 1(a), we fit a linear regression model using the logarithmic transform of $C,S$ to predict the tpr (also logarithmically transformed) with statsmodels Seabold & Perktold (2010). The general form of the model can be found in Equation 3 where $\beta_{S},\beta_{C}$ and $\beta_{0}$ are the learnable regression parameters.
$\log_{10}(\textsc{tpr})=\beta_{S}\log_{10}(S)+\beta_{C}\log_{10}(C)+\beta_{0}$ (3)
We utilize MIA results of ViT-B (see Table A3) as the training data and use R-50 (see Table A4) as the test data to investigate if the vulnerability prediction model generalizes to a completely different backbone.
The parameters of the prediction model fitted to the training data (note that the parameters are specific to $\textsc{fpr}=0.001$) are $\beta_{S}=-0.492,\beta_{C}=0.357$ and $\beta_{0}=-0.486$. Based on the $R^{2}$ score ($R^{2}=0.917$), our model fits the data extremely well. Furthermore, Figure 3 shows that the model is robust to a change of feature extractor, as it is able to predict the tpr for R-50 despite being trained on ViT-B (test $R^{2}=0.740$). See Section C.1.3 for further supporting results at $\textsc{fpr}\in\{0.1,0.01\}$.
Figure 3: Predicted MIA vulnerability (tpr at $\textsc{fpr}=0.001$) as a function of $S$ (shots) using a model based on Equation 3 fitted Table A3 (ViT-B). The dots show the median tpr for the train set (ViT-B; Table A3) and the test set (R-50; Table A4) over six seeds (datasets: Patch Camelyon, EuroSAT and CIFAR100).
6 MIA vulnerability for data samples
This Section extends our investigation from the data set level properties studied in Section 5 to individual data sample level properties. Previously Yu et al. (2023) have studied correlations between classes and privacy risk before, but in comparison to their study, we look at data samples independent from classes.
We will focus on a fine-tuned last layer classifier (Head) trained on top of a ViT-B, pre-trained on ImageNet-21k Russakovsky et al. (2015). We investigate two data sample properties that could potentially influence the vulnerability (tpr at fpr) of data points:
  • Per-sample gradients: Per-sample gradients form the basis for DP individual accounting Feldman & Zrnic (2021); Koskela et al. (2023). In individual accounting a larger sum of gradient norms results in a higher individual privacy loss. Our hypothesis is that this holds also for MIA vulnerability in non-DP settings.
  • Vector space distances: The pre-trained feature extractors project images into a vector space in which one can compute distances between individual data samples. Our hypothesis is that data samples that are further away from other data samples are easier to distinguish.
We study individual samples, but aggregate the results over multiple data samples by binning the samples based on their data sample properties to reduce the noise. For each data sample we compute its property and assign the sample to a bin based on its percentile in relation to other samples. We then compute the tpr at fpr for each bin. We finally report for each data property bin the percentile of the tpr at fpr.
Figure 4: MIA vulnerability (tpr at fpr; percentiles among a seed) as a function of data sample properties (binned among a seed). The results have been obtained over 6 seeds with $S=256$ (Patch Camelyon) and $S=64$ (CIFAR10).
Figure 4 shows the observed relationships between the MIA vulnerability and the average individual gradient norms and the vector space distance. We find that:
  • Per-sample gradients: The average gradient norms are correlated with the vulnerability in all our experiments, but the strength of the correlation depends how many steps are taken into account. Generally, we observe that the correlation is stronger when we only consider the gradients at the end of the training (either the last 20 or 5 steps). This is different from the individual accounting methods that rely on the whole gradient trace for DP accounting.
  • Vector space distances: The distance of data samples to their nearest neighbor correlates with the vulnerability for CIFAR10, but not for Patch Camelyon. We leave a more comprehensive study of this correlation and its limitations for future work.
7 Discussion
As DP reduces the utility of models, it is important to understand also when it is necessary. Our results on MIA vulnerability as a function of data set properties display a strong power law dependence on the number of examples per class. This leads to an unfortunate dilemma: DP seems most important in the small data regime where it likely reduces the model utility the most.
Limitations
Due to computational constraints, we only report results on transfer learning with fine-tuning only a part of the model.
Acknowledgements
This work was supported by the Research Council of Finland (Flagship programme: Finnish Center for Artificial Intelligence, FCAI, Grant 356499 and Grant 359111), the Strategic Research Council at the Research Council of Finland (Grant 358247) as well as the European Union (Project 101070617). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the granting authority can be held responsible for them. The authors wish to thank the Finnish Computing Competence Infrastructure (FCCI) for supporting this project with computational and data storage resources. We thank John F. Bronskill for helpful discussions regarding few-shot learning.
Appendix A Background
Differential privacy (DP; Dwork et al., 2006b) is the gold standard for formalizing privacy guarantees. $(\epsilon,\delta)$-DP (Dwork et al., 2006a) has a privacy budget consisting of $\epsilon\geq 0$ and $\delta\in[0,1]$, where smaller values correspond to a stronger privacy guarantee. We refer to Dwork & Roth (2014) for a comprehensive intro to DP.
DP deep learning utilizes DP-SGD (Rajkumar & Agarwal, 2012; Song et al., 2013; Abadi et al., 2016), which is the DP adaptation of stochastic gradient descent. DP-SGD clips per-sample gradients to a pre-determined norm to limit the influence of each sample and adds Gaussian noise the sum of the clipped gradients.
Clipping the gradients and adding Gaussian noise in DP-SGD typically comes with a cost in accuracy. This decrease in accuracy can be alleviated through transfer-learning. Publically available pre-trained models are fine-tuned with DP-SGD on sensitive downstream data sets. Both in NLP (Li et al., 2022; Yu et al., 2022) and computer vision (Kurakin et al., 2022; De et al., 2022; Mehta et al., 2022; Cattan et al., 2022; Tobaben et al., 2023) transfer-learning significantly closes the gap between non-private and private training. Tramèr et al. (2022) critically discuss this line of work and the implication of relying on pre-trained models.
Appendix B Training details
B.1 Parameterization
We utilize pre-trained feature extractors BiT-M-R50x1 (R-50) (Kolesnikov et al., 2020) with 23.5M parameters and Vision Transformer ViT-Base-16 (ViT-B) (Dosovitskiy et al., 2021) with 85.8M parameters, both pretrained on the ImageNet-21K data set (Russakovsky et al., 2015). We download the feature extractor checkpoints from the respective repositories.
Following Tobaben et al. (2023) that show the favorable trade-off of parameter-efficient fine-tuning between computational cost, utility and privacy even for small data sets, we only consider fine-tuning subsets of all feature extractor parameters. We consider the following configurations:
  • Head: We train a linear layer on top of the pre-trained feature extractor.
  • FiLM: In addition to the linear layer from Head, we fine-tune parameter-efficient FiLM (Perez et al., 2018) adapters scattered throughout the network. While a diverse set of adapters has been proposed, we utilize FiLM as it has been shown to be competitive in prior work (Shysheya et al., 2023; Tobaben et al., 2023).
B.2 Hyperparameter tuning
Our hyperparameter tuning is heavily inspired by the comprehensive few-shot experiments by Tobaben et al. (2023). We utilize their hyperparameter tuning protocol as it has been proven to yield SOTA results for (DP) few-shot models.
[…]
Table A1: Hyperparameter ranges used for the Bayesian optimization with Optuna.
lower bound upper bound
batch size 10 $|\mathcal{D}|$
clipping norm 0.2 10
epochs 1 200
learning rate 1e-7 1e-2
noise multiplier Based on target $\epsilon$
B.3 Datasets
Table A2 shows the used data sets in the paper. We base our experiments on a subset of the the few-shot benchmark VTAB (Zhai et al., 2019) that achieves a classification accuracy $>80\%$ and thus would considered to be used by a practitioner. Additionally, we add CIFAR10 that is not part of the original VTAB benchmark.
Table A2: Used datasets in the paper, their minimum and maximum shots $S$ and maximum number of classes $C$ and their test accuracy when fine-tuning a non-DP ViT-B Head. The test accuracy for EuroSAT and Resics45 is computed on the part of the training split that is not used for training the particular model due to both datasets missing an official test split. Note that LiRA requires $2S$ for training the shadow models and thus $S$ is smaller than when only performing fine-tuning.
dataset (max.) min. max. test accuracy test accuracy
$C$ $S$ $S$ (min. $S$) (max. $S$)
Patch Camelyon (Veeling et al., 2018) 2 256 65536 82.8% 85.6%
CIFAR10 (Krizhevsky, 2009) 10 8 2048 92.7% 97.7%
EuroSAT (Helber et al., 2019) 10 8 512 80.2% 96.7%
Pets (Parkhi et al., 2012) 37 8 32 82.3% 90.7%
Resics45 (Cheng et al., 2017) 45 32 256 83.5% 91.6%
CIFAR100 (Krizhevsky, 2009) 100 16 128 82.2% 87.6%
Appendix C Additional results
In this section, we provide tabular results for our experiments and additional figures that did not fit into the main paper.
(a) tpr at $\textsc{fpr}=0.1$
(b) tpr at $\textsc{fpr}=0.01$
(c) tpr at $\textsc{fpr}=0.001$
Figure A.1: MIA vulnerability as a function of privacy budget ($\epsilon$ at $\delta=10^{-5}$) when attacking different backbones fine-tuned on CIFAR10 and CIFAR100 at $S\in\{25,50\}$. The solid line displays the median and the error bars the minimum of the lower bounds and the maximum of the upper bounds for the Clopper-Pearson CIs over 15 seeds. The DP bounds are the theoretical DP bounds based on Kairouz et al. (2015) where the naive DP bound is computed with $\delta=10^{-5}$ and the tight DP bound over an interval of deltas and their respective $\epsilon$.
C.1 Additional results for Section 5
This section contains additional results for Section 5.
C.1.1 Vulnerability as a function of shots
This section displays additional results to Figure 1(a) for fpr $\in\{0.1,0.01,0.001\}$ for ViT-B and R-50 in in Figures A.2, A3 and A4.
(a) ViT-B Head tpr at $\textsc{fpr}=0.1$
(b) R-50 Head tpr at $\textsc{fpr}=0.1$
(c) ViT-B Head tpr at $\textsc{fpr}=0.01$
(d) R-50 Head tpr at $\textsc{fpr}=0.01$
(e) ViT-B Head tpr at $\textsc{fpr}=0.001$
(f) R-50 Head tpr at $\textsc{fpr}=0.001$
Figure A.2: MIA vulnerability as a function of shots (examples per class) when attacking a pre-trained ViT-B and R-50 Head trained without DP on different downstream data sets. The errorbars display the min and max Clopper-Pearson CIs over 6 seeds and the solid line the median.
Table A3: Median MIA vulnerability over six seeds as a function of $S$ (shots) when attacking a Head trained without DP on-top of a ViT-B. The ViT-B is pre-trained on ImageNet-21k.
dataset classes ($C$) shots ($S$) tpr@fpr=0.1 tpr@fpr=0.01 tpr@fpr=0.001 Patch Camelyon (Veeling et al., 2018) 2 256 0.266 0.086 0.032 512 0.223 0.059 0.018 1024 0.191 0.050 0.015 2048 0.164 0.037 0.009 4096 0.144 0.028 0.007 8192 0.128 0.021 0.005 16384 0.118 0.017 0.003 32768 0.109 0.014 0.002 65536 0.105 0.012 0.002 CIFAR10 (Krizhevsky, 2009) 10 8 0.910 0.660 0.460 16 0.717 0.367 0.201 32 0.619 0.306 0.137 64 0.345 0.132 0.067 128 0.322 0.151 0.082 256 0.227 0.096 0.054 512 0.190 0.068 0.032 1024 0.168 0.056 0.025 2048 0.148 0.039 0.013 EuroSAT (Helber et al., 2019) 10 8 0.921 0.609 0.408 16 0.738 0.420 0.234 32 0.475 0.222 0.113 64 0.400 0.159 0.074 128 0.331 0.155 0.084 256 0.259 0.104 0.049 512 0.213 0.080 0.037 Pets (Parkhi et al., 2012) 37 8 0.648 0.343 0.160 16 0.745 0.439 0.259 32 0.599 0.311 0.150 Resics45 (Cheng et al., 2017) 45 32 0.672 0.425 0.267 64 0.531 0.295 0.168 128 0.419 0.212 0.115 256 0.323 0.146 0.072 CIFAR100 (Krizhevsky, 2009) 100 16 0.814 0.508 0.324 32 0.683 0.445 0.290 64 0.538 0.302 0.193 128 0.433 0.208 0.114
Table A4: Median MIA vulnerability over six seeds as a function of $S$ (shots) when attacking a Head trained without DP on-top of a R-50. The R-50 is pre-trained on ImageNet-21k.
dataset classes ($C$) shots ($S$) tpr@fpr=0.1 tpr@fpr=0.01 tpr@fpr=0.001 Patch Camelyon (Veeling et al., 2018) 2 256 0.272 0.076 0.022 512 0.195 0.045 0.011 1024 0.201 0.048 0.011 2048 0.178 0.041 0.010 4096 0.163 0.033 0.008 8192 0.143 0.026 0.006 16384 0.124 0.019 0.004 32768 0.118 0.016 0.003 65536 0.106 0.012 0.002 CIFAR10 (Krizhevsky, 2009) 10 8 0.911 0.574 0.324 16 0.844 0.526 0.312 32 0.617 0.334 0.183 64 0.444 0.208 0.106 128 0.334 0.159 0.084 256 0.313 0.154 0.086 512 0.251 0.103 0.051 1024 0.214 0.082 0.038 EuroSAT (Helber et al., 2019) 10 8 0.846 0.517 0.275 16 0.699 0.408 0.250 32 0.490 0.236 0.121 64 0.410 0.198 0.105 128 0.332 0.151 0.075 256 0.269 0.111 0.056 512 0.208 0.077 0.036 Pets (Parkhi et al., 2012) 37 8 0.937 0.631 0.366 16 0.745 0.427 0.227 32 0.588 0.321 0.173 Resics45 (Cheng et al., 2017) 45 32 0.671 0.405 0.235 64 0.534 0.289 0.155 128 0.445 0.231 0.121 256 0.367 0.177 0.088 CIFAR100 (Krizhevsky, 2009) 100 16 0.897 0.638 0.429 32 0.763 0.549 0.384 64 0.634 0.414 0.269
C.1.2 Vulnerability as a function of the number of classes
This section displays additional results to Figure 1(b) for fpr $\in\{0.1,0.01,0.001\}$ for ViT-B and R-50 in in Figures A.3, A5 and A6.
(a) ViT-B Head tpr at $\textsc{fpr}=0.1$
(b) R-50 Head tpr at $\textsc{fpr}=0.1$
(c) ViT-B Head tpr at $\textsc{fpr}=0.01$
(d) R-50 Head tpr at $\textsc{fpr}=0.01$
(e) ViT-B Head tpr at $\textsc{fpr}=0.001$
(f) R-50 Head tpr at $\textsc{fpr}=0.001$
Figure A.3: MIA vulnerability as a function of $C$ (classes) when attacking a ViT-B and R-50 Head fine-tuned without DP on different data sets where the classes are randomly sub-sampled and $S=32$. The solid line displays the median and the errorbars the min and max clopper-pearson CIs over 12 seeds.
Table A5: Median MIA vulnerability over 12 seeds as a function of $C$ (classes) when attacking a Head trained without DP on-top of a ViT-B. The Vit-B is pre-trained on ImageNet-21k.
dataset shots ($S$) classes ($C$) tpr@fpr=0.1 tpr@fpr=0.01 tpr@fpr=0.001 Patch Camelyon (Veeling et al., 2018) 32 2 0.467 0.192 0.080 CIFAR10 (Krizhevsky, 2009) 32 2 0.494 0.167 0.071 4 0.527 0.217 0.115 8 0.574 0.262 0.123 EuroSAT (Helber et al., 2019) 32 2 0.306 0.100 0.039 4 0.298 0.111 0.047 8 0.468 0.211 0.103 Pets (Parkhi et al., 2012) 32 2 0.232 0.045 0.007 4 0.324 0.092 0.035 8 0.296 0.094 0.035 16 0.406 0.158 0.069 32 0.553 0.269 0.136 Resics45 (Cheng et al., 2017) 32 2 0.272 0.084 0.043 4 0.322 0.119 0.056 8 0.496 0.253 0.148 16 0.456 0.204 0.108 32 0.580 0.332 0.195 CIFAR100 (Krizhevsky, 2009) 32 2 0.334 0.088 0.035 4 0.445 0.150 0.061 8 0.491 0.223 0.121 16 0.525 0.256 0.118 32 0.553 0.276 0.153 64 0.612 0.350 0.211
Table A6: Median MIA vulnerability over 12 seeds as a function of $C$ (classes) when attacking a Head trained without DP on-top of a R-50. The R-50 is pre-trained on ImageNet-21k.
dataset shots ($S$) classes ($C$) tpr@fpr=0.1 tpr@fpr=0.01 tpr@fpr=0.001 Patch Camelyon (Veeling et al., 2018) 32 2 0.452 0.151 0.041 CIFAR10 (Krizhevsky, 2009) 32 2 0.404 0.146 0.060 4 0.560 0.266 0.123 8 0.591 0.318 0.187 EuroSAT (Helber et al., 2019) 32 2 0.309 0.111 0.050 4 0.356 0.144 0.064 8 0.480 0.233 0.123 Pets (Parkhi et al., 2012) 32 2 0.249 0.068 0.029 4 0.326 0.115 0.056 8 0.419 0.173 0.075 16 0.493 0.245 0.127 32 0.559 0.294 0.166 Resics45 (Cheng et al., 2017) 32 2 0.310 0.103 0.059 4 0.415 0.170 0.083 8 0.510 0.236 0.119 16 0.585 0.311 0.174 32 0.644 0.382 0.218 CIFAR100 (Krizhevsky, 2009) 32 2 0.356 0.132 0.054 4 0.423 0.176 0.087 8 0.545 0.288 0.163 16 0.580 0.338 0.196 32 0.648 0.402 0.244 64 0.711 0.476 0.320
C.1.3 Predicting data set vulnerability as function of $S$ and $C$
Table A7: Results for fitting Equation 3 with statsmodels Seabold & Perktold (2010) to ViT Head data at fpr $\in\{0.1,0.01,0.001\}$. We utilize an ordinary least squares. The test $R^{2}$ assesses the fit to the data of R-50 Head.
coeff. fpr $R^{2}$ test $R^{2}$ coeff. value std. error $t$ $p>|z|$ coeff. $[0.025$ coeff. $0.975]$ $\beta_{S}$ (for $S$) 0.1 0.924 0.898 -0.237 0.008 -30.994 $0$0.252 -0.222 0.01 0.937 0.839 -0.380 0.012 -31.818 $0$0.403 -0.356 0.001 0.917 0.740 -0.492 0.019 -26.143 $0$0.530 -0.455 $\beta_{C}$ (for $C$) 0.1 0.924 0.898 0.095 0.014 6.992 0.000 0.068 0.121 0.01 0.937 0.839 0.226 0.021 10.680 0.000 0.184 0.268 0.001 0.917 0.740 0.357 0.033 10.677 0.000 0.291 0.423 $\beta_{0}$ (intercept) 0.1 0.924 0.898 -0.008 0.029 -0.274 0.785 -0.064 0.049 0.01 0.937 0.839 -0.224 0.045 -5.006 $0$0.312 -0.136 0.001 0.917 0.740 -0.486 0.071 -6.879 $0$0.626 -0.347
(a) tpr at $\textsc{fpr}=0.1$
(b) tpr at $\textsc{fpr}=0.01$
(c) tpr at $\textsc{fpr}=0.001$
Figure A.4: Predicted MIA vulnerability as a function of $S$ (shots) using a model based on Equation 3 fitted Table A3 (ViT-B). The dots show the median tpr for the train set (ViT-B; Table A3) and the test set (R-50; Table A4) over six seeds. Note that the dots for $C=10$ are for EuroSAT.