### v2

# On the Impact of Dataset Properties on Membership Privacy of Deep Learning

Abstract
We apply a state-of-the-art membership inference attack (MIA) to systematically test the practical privacy vulnerability of fine-tuning large image classification models. We focus on understanding the properties of data sets and samples that make them vulnerable to membership inference. In terms of data set properties, we find a strong power law dependence between the number of examples per class in the data and the MIA vulnerability, as measured by true positive rate of the attack at a low false positive rate. We train a linear model to predict true positive rate based on data set properties and observe good fit for MIA vulnerability on unseen data. To analyse the phenomenon theoretically, we reproduce the result on a simplified model of membership inference that behaves similarly to our experimental data. We prove that in this model, the logarithm of the difference of true and false positive rates depends linearly on the logarithm of the number of examples per class. For an individual sample, the gradient norm is predictive of its vulnerability.
1 Introduction
Machine learning models are prone to memorising their training data, which makes them vulnerable to privacy attacks such as membership inference attacks (MIAs; Shokri et al., 2017; Carlini et al., 2022) and reconstruction attacks (e.g. Balle et al., 2022; Nasr et al., 2023). Differential privacy (DP; Dwork et al., 2006) provides protection against these attacks, but strong formal protection often comes at the cost of significant loss of model utility.
Finding the correct balance between making models resistant to attacks while maintaining a high utility is important for many applications. In health, for example, many European countries and soon also the EU within the European Health Data Space have requirements that models trained on health data that are made publicly available must be anonymous, i.e. they must not contain information that can be linked to an identifiable individual. On the other hand, loss of utility of the model due to privacy constraints may compromise the health benefits that might be gained from it.
In this paper, our aim is to apply a state-of-the-art MIA systematically, to help understand practical privacy risks when training deep-learning-based classifiers without DP protections. Our case study focuses on understanding and quantifying factors that influence the vulnerability of non-DP deep learning models to MIA. Since the earliest MIA methods, there has been evidence that classification models with more classes are more vulnerable to MIA (Shokri et al., 2017). Similarly, there is evidence that models trained on fewer samples can be more vulnerable (Chen et al., 2020; Németh et al., 2023). Based on extensive experimental data over many data sets with varying sizes, we uncover a power law describing the vulnerability remarkably well.
Previously, Tobaben et al. (2023) and Yu et al. (2023) have presented limited studies of how data set properties affect the MIA vulnerability. Tobaben et al. (2023) reported how the MIA vulnerability of few-shot image classification is affected by the number of shots (i.e. the number of examples available per class). Yu et al. (2023) studied how the MIA vulnerability of a class in an image classifier trained from scratch depends on the average individual privacy parameters of that class. Our work significantly expands on both of these by more systematic study and modelling of data set properties and by providing a more detailed look at individual samples, not just classes.
List of contributions In this work we perform an extensive empirical study on the MIA vulnerability of deep learning models. We focus on a transfer learning setting for image classification tasks, where a large pre-trained neural network is fine-tuned on a sensitive data set, and the MIA is performed on the final classifier. We find that the MIA vulnerability has a strong correlation with certain data set properties (Section 3). In Section 4 we analyse the reasons for these findings more thoroughly and we extend the analysis to individual samples in Section 5. The primary contributions of this work are:
  1. 1.
    Few-shot MIA: We conduct a comprehensive study of MIA vulnerability (tpr at fixed low fpr) of target models trained using many different data sets with varying sizes and uncover a power law between the number of examples per class and the vulnerability to MIA (see Figure 1).
  2. 2.
    Regression model: We utilise our empirical observations to train a regression model to predict MIA vulnerability ($\log(\textsc{tpr}-\textsc{fpr})$ at fixed low fpr) based on examples per class ($\log S$) and number of classes ($\log C$). We show both very good fit on the training data as well as good prediction quality on unseen data from a different feature extractor and when fine-tuning other parameterisations (see Figure 3).
  3. 3.
    Theoretical analysis using a simplified model of MIA: To analyse the phenomenon theoretically, we reproduce the result on a simplified model of membership inference that behaves similarly to our experimental data. In this model we can formally prove that $\log(\textsc{tpr}-\textsc{fpr})$ depends linearly on $\log S$ (see Section 4).
  4. 4.
    Individual sample MIA: We extend the study to individual samples and find: 1 The gradient norm is predictive of vulnerability of an individual sample. 2 Gradient norms at different stages of learning have different impact, but how depends on the data set (see Section 5).
2 Background and Methods
Notation for the properties of the training data set $\mathcal{\mathcal{D}}$: 1 $C$for the number of classes 2 $S$for shots (examples per class) 3 $\mathcal{|\mathcal{D}|}$for training data set size ($\mathcal{|\mathcal{D}|}=CS$). We denote the number of MIA shadow models with $M$.
Membership inference attacks (MIAs) aim to infer whether a particular sample was part of the training set of the targeted model Shokri et al. (2017). Thus, they can be used to determine lower bounds on the privacy leakage of models to complement the theoretical upper bounds obtained through differential privacy.
Likelihood Ratio attack (LiRA; Carlini et al., 2022) While many different MIAs have been proposed Hu et al. (2022), in this work we consider the SOTA: the Likelihood Ratio Attack (LiRA). LiRA assumes an attacker that has black-box access to the attacked model, knows the training data distribution, the training set size, the model architecture, hyperparameters and training algorithm. Based on this information, the attacker can train so-called shadow models (Shokri et al., 2017) which imitate the model under attack but for which the attacker knows the training data set.
LiRA exploits the observation that the loss function value used to train a model is often lower for examples that were part of the training set compared to those that were not. For a target sample $x$, LiRA trains the shadow models: 1 with $x$ as a part of the training set ($x\in\mathcal{D}$) 2 and without $x$ in the training set ($x\notin\mathcal{D}$) . After training the shadow models, $x$ is passed through the shadow models, and based on the losses (or predictions) two Gaussian distributions are formed: one for the losses of $x\in\mathcal{D}$ shadow models, and one for the $x\notin\mathcal{D}$. Finally, the attacker computes the loss for the point $x$ using the model under attack and determines using a likelihood ratio test on the distributions built from the shadow models whether it is more likely that $x\in\mathcal{D}$ or $x\notin\mathcal{D}$. Carlini et al. (2022) proposed an optimisation for performing LiRA for multiple models and points without training a computationally infeasible number of shadow models. We utilise this optimisation.
Measuring MIA vulnerability Using the likelihood-ratios from LiRA as a score, we can build a binary classifier to predict whether a sample belongs to the training data or not. The operational characteristics of such classifier can be used to measure the success of the MIA. More specifically, throughout the rest of the paper, we will use the true positive rate (tpr) at a specific false positive rate (fpr) as a measure for the vulnerability. Identifying even a small number of samples with high confidence is considered harmful (Carlini et al., 2022) and thus we focus on the regions of small fpr.
Measuring the uncertainty for tpr The tpr values from the LiRA-based classifier can be seen as maximum likelihood-estimators for the probability of producing true positives among the positive samples. Since we have a finite number of samples for our estimation, it is important to estimate the uncertainty in these estimators. Therefore, when we report the tpr values for a single repeat of the learning algorithm, we estimate the stochasticity of the tpr estimate by using Clopper-Pearson intervals Clopper and Pearson (1934). Given tp true positives among p positives, the $1-\alpha$ confidence Clopper-Pearson interval for the tpr is given as
$\begin{split}&B(\alpha/2;\textsc{tp},\textsc{p}-\textsc{tp}+1)<\textsc{tpr}\\ &\textsc{tpr}<B(1-\alpha/2;\textsc{tp}+1,\textsc{p}-\textsc{tp}),\end{split}$ (1)
where $B(q;a,b)$ is the $q$th-quantile of $\text{Beta}(a,b)$ distribution.
3 MIA Vulnerability and Data Set Properties
In this section, we investigate how different properties of data sets affect the MIA vulnerability. Based on our observations, we propose a method to predict the vulnerability to MIA using these properties.
3.1 Experimental Setup
We focus on a image classification setting where we fine-tune pre-trained models on sensitive downstream data sets and assess the MIA vulnerability using LiRA with $M=256$ shadow models. We base our experiments on a subset of the few-shot benchmark VTAB (Zhai et al., 2019) that achieves a test classification accuracy $>80\%$ (see Table A2).
We report results for fine-tuning a last layer classifier (Head) trained on top of a Vision Transformer ViT-Base-16 (ViT-B; Dosovitskiy et al., 2021), pre-trained on ImageNet-21k Russakovsky et al. (2015). The results for using ResNet-50 (R-50; Kolesnikov et al., 2020) as a backbone can be found in Section B.1. We optimise the hyperparameters (batch size, learning rate and number of epochs) using the library Optuna Akiba et al. (2019) with the Tree-structured Parzen Estimator (TPE; Bergstra et al., 2011) sampler with 20 iterations (more details in Section A.2). We provide the the code for reproducing the experiments in the supplementary material.
3.2 Experimental Results
Using the setting described above, we study how the number of classes and the number of shots affect the vulnerability (tpr at fpr as described in Section 2). We make the following observations:
  • A larger number of $S$ (shots) decrease the vulnerability in a power law relation as demonstrated in Figure 1(a). We provide further evidence of this in the Appendix (Figures A.1, A3 and A4).
  • Contrary, a larger number of $C$ (classes) increases the vulnerability as demonstrated in Figure 1(b) with further evidence in Figures A.2, A5 and A6 in the Appendix. However, the trend w.r.t. $C$ is not as clear as with $S$.
(a) as function of $S$ (shots)
(b) as function of $C$ (classes) at $S=32$
Figure 1: MIA vulnerability (tpr at $\textsc{fpr}=0.001$) as a function of data set properities when attacking a ViT-B Head fine-tuned without DP on different data sets. The solid line displays the median and the error bars the minimum of the lower bounds and maximum of the upper bounds for the Clopper-Pearson CIs over multiple seeds (six for Figure 1(a) and 12 for Figure 1(b))
3.3 Model to Predict Data Set Vulnerability
[…]
$\log_{10}(\textsc{tpr}-\textsc{fpr})=\beta_{S}\log_{10}(S)+\beta_{C}\log_{10}(C)+\beta_{0}$ (2)
[…]
We utisee MIA results of ViT-B (see Table A3) as the training data. Based on the $R^{2}$ (coefficient of determination) score ($R^{2}=0.930$ for the model trained on $\textsc{fpr}=0.001$ data), our model fits the data extremely well. We provide further evidence for other fpr in Figure A.3 and Table A8 in the Appendix. Figure 2 shows the parameters of the prediction model fitted to the training data.
Figure 2: Coefficient values for different fpr when fitting a regression model based on Equation 2 fitted on data from ViT-B (Head) (Table A3). The error bars display the 95% confidence intervals based on Student’s t-distribution.
Prediction quality on other MIA target models
We analyse how the regression model trained on the ViT-B (Head) data generalizes to other target models. The main points are:
  • R-50 (Head): Figure 3(a) shows that the regression model is robust to a change of the feature extractor, as it is able to predict the tpr for R-50 (Head) (test $R^{2}=0.790$).
  • R-50 (FiLM): Figure 3(b) shows that the prediction quality is good for R-50 (FiLM) models. These models are fine-tuned with parameter-efficient FiLM (Perez et al., 2018) layers (See Section A.1). Tobaben et al. (2023) demonstrated that FiLM layers are a competitive alternative to training all parameters. We supplement the MIA results of Tobaben et al. (2023) with own FiLM training runs. Refer to Table A7 in the Appendix.
  • From-Scratch-Training: Carlini et al. (2022) provide limited results on from-scratch-training. To the best of our knowledge these are the only published LiRA results on image classification models. Figure 3(b) displays that our prediction model underestimates the vulnerability of the from-scratch trained target models. We have identified two potential explanations for this 1 In from-scratch-training all weights of the model need to be trained from the sensitive data and thus potentially from-scratch-training could be more vulnerable than fine-tuning. 2 The strongest attack in Carlini et al. (2022) uses data augmentations to improve the performance. We are not using this optimization.
[…]
(b) Evaluating prediction performance for target models 1 left: fine-tuned with FiLM (data partially from Tobaben et al. (2023), see Table A7) 2 right: trained from scratch (data from Carlini et al. (2022))
Figure 3: Performance of the regression model based on Equation 2 fitted on data from Table A3.
4 Understanding the MIA Prediction Model
In this section, we focus on understanding why the previously observed trends exist. Our goal is to formally prove a similar relationship as in the predictive model of Section 3.3.
4.1 A Simplified Model of Membership Inference
In this section, we construct a simplified model of membership inference that streamlines the data generation and shadow model training. We observe that the trends of the vulnerability of the simplified model is a good approximation of the trends previously observed with the image classification example.
Instead of training linear classifiers to classify images, we sample vectors on a high dimensional unit sphere and classify based on inner product with estimated class mean. This model is easier to analyse theoretically, and allows us to simulate arbitrary combinations of $S$ and $C$ without relying on the existence of a specific image classification downstream data set as especially high $C$ image classification downstream data sets that are computationally feasible to fine-tune on have small $S$. We generate the data and form the classifiers (which are our target models) as follows:
  1. 1.
    For each class, we first sample a true class mean $m_{c}$ on a high dimensional unit sphere that is orthogonal to all other true class means ($\forall i,j\in C:m_{i}\perp m_{j}\lor i=j$).
  2. 2.
    We sample $2S$ vectors $x_{c}$ for each class. We assume that they are Gaussian distributed around the the true class mean $x_{c}\sim\mathcal{N}(m_{c},s^{2}I)$ where the $s^{2}$ is the in-class variance.
  3. 3.
    For each "target model" we randomly choose a subset of size $CS$ from all generated vectors and compute per-class means $r_{c}$.
  4. 4.
    The computed mean is used to classify sample $(x,y)$ by computing the inner product $\langle x,r_{c}\rangle$ as a metric of similarity.
The attacker has to infer which vectors have been used for training the classifier. Instead of utilising the logits (like in the image classification example), the attacker can use the inner products of a point with the cluster means. In Figure 4 we compare the attacker’s performance as a function of $C$ and $S$ to the the previous image classification example.
This simplified model resembles the linear (Head) classifier. In the linear classifier, we find a matrix $A$ and biases $b$, to optimize the cross-entropy between the labels and logits $Az+b$, where $z$ denotes the feature space representation of the data. In the simplified model, the rows of $A$ are replaced by the cluster means and we do not include the bias term in the classification.
(a) as a function of $S$ (shots)
(b) as a function of $C$ (classes)
[…]
4.2 Theoretical Analysis
Similarly to the setting in Section 4.1, we focus on a classification model that aims to classify the sample $(x,y)$ based on the inner product $\langle x,r_{c}\rangle$, where $r_{c}$ is the mean of the training data points $x_{m}$ with $y_{m}=c$. However, we only focus on the scores of the target class, i.e. $c=y$, and therefore the score used of LiRA simplifies to $\langle x,r_{y}\rangle$.
The fpr for LiRA is given as
$\textsc{fpr}=\Pr\left(\frac{\mathcal{N}(t_{x};\mu_{\textrm{in}},\sigma_{\textrm{in}}^{2})}{\mathcal{N}(t_{x};\mu_{\textrm{out}},\sigma_{\textrm{out}}^{2})}>\exp(\tau)\mid x\notin\mathcal{D}_{\textrm{target}}\right),$ (3)
where $t_{x}=\langle x,r_{y}^{\textrm{(target)}}\rangle$ is the score for the target model, $\tau$ is the decision threshold for the classifier, and $\mu_{\textrm{in}},\mu_{\textrm{out}},\sigma_{\textrm{in}}$ and $\sigma_{\textrm{out}}$ are the LiRA statistics computed from the shadow models. We assume that $\sigma_{\textrm{in}}=\sigma_{\textrm{out}}=\sigma$, which allows to write Equation 3 as
fpr $=\Pr\left(\log\frac{\mathcal{N}(t_{x};\mu_{\textrm{in}},\sigma^{2})}{\mathcal{N}(t_{x};\mu_{\textrm{out}},\sigma^{2})}>\tau\mid x\notin\mathcal{D}_{\textrm{target}}\right)$ (4)
$=\Pr\left(-\frac{1}{2\sigma^{2}}(\mu_{\textrm{in}}-\mu_{\textrm{out}})\left[-2t_{x}+(\mu_{\textrm{in}}+\mu_{\textrm{out}})\right]>\tau\mid x\notin\mathcal{D}_{\textrm{target}}\right).$ (5)
From now on, we will denote the $-\frac{1}{2\sigma^{2}}(\mu_{\textrm{in}}-\mu_{\textrm{out}})\left[-2t_{x}+(\mu_{\textrm{in}}+\mu_{\textrm{out}})\right]$ with $\lambda_{\textsc{fpr}}$ when $x\notin\mathcal{D}_{\textrm{target}}$ and with $\lambda_{\textsc{tpr}}$ otherwise. Furthermore we will drop the condition $x\in\mathcal{D}_{\textrm{target}}$ or $x\notin\mathcal{D}_{\textrm{target}}$ from the fpr and tpr expressions
fpr $:=\Pr(\lambda_{\textsc{fpr}}>\tau)$ (6)
tpr $:=\Pr(\lambda_{\textsc{tpr}}>\tau).$ (7)
Lemma 1.
In the limit of infinite shadow models and with large enough $S$, the LiRA test statistics $\lambda_{\textsc{fpr}}$ and $\lambda_{\textsc{tpr}}$ obtain the following forms
$\lambda_{\textsc{fpr}}$ $=-\frac{1}{2s^{2}||x||^{2}}\langle x,x-m_{y}\rangle\left(\frac{2s}{\sqrt{S}}\langle x,z\rangle+\frac{1}{S}\langle x,x-m_{y}\rangle\right)$ (8)
$\lambda_{\textsc{tpr}}$ $\approx-\frac{1}{2s^{2}||x||^{2}}\langle x,x-m_{y}\rangle\left(\frac{2s}{\sqrt{S}}\langle x,z\rangle-\frac{2}{S}||x||^{2}+\frac{1}{S}\langle x,x-m_{y}\rangle\right),$ (9)
where $z\sim\mathcal{N}(0,I_{d})$.
Proof.
We defer the proof to Appendix C. ∎
The factors independent of $S$ in for both $\lambda_{\textsc{fpr}}$ and $\lambda_{\textsc{tpr}}$ correspond to the $-(\mu_{\textrm{in}}-\mu_{\textrm{out}})/(2\sigma^{2})$. Figure A.5 illustrates how this factor behaves as a function of $S$ in our experiments with the simple model. As predicted by the theory, the factor (i.e. the difference in means divided by the variance) remains fairly constant across different values of $S$.
Using Lemma 1 we can derive a rate for the decision threshold $\tau$:
Corollary 2.
In order to maintain constant fpr, we need to adjust $\tau$ as $O(1/\sqrt{S})$.
According to Corollary 2, the $\tau:=\tau(S)$ evolves as $O(1/\sqrt{S})$, and we will approximate it as $\tau(S)=\gamma_{1}\frac{1}{\sqrt{S}}$, with constant $\gamma_{1}$ depending on the target fpr.
Theorem 3.
$\log(\textsc{tpr}-\textsc{fpr})$ depends linearly on $\log(S)$.
Proof.
From Lemma 1, we see that $\lambda_{\textsc{fpr}}$ and $\lambda_{\textsc{tpr}}$ differ only in a single term which evolves as $O(1/S)$. Thus for sufficiently large $S$ we can consider this additional r.v. as a small constant shift to the distribution of $\lambda_{\textsc{fpr}}$. Let us denote this variable now with $q$. When $S$ grows, $q$ becomes small, and we have
[…]
$\approx p_{\lambda_{\textsc{fpr}}}(\tau)q.$ (11)
Taking the log yields
$\log(\textsc{tpr}-\textsc{fpr})$ $\approx\log\left(p_{\lambda_{\textsc{fpr}}}\left(\tau\right)\right)+\log q$ (12)
$=\log\left(p_{\lambda_{\textsc{fpr}}}\left(\tau\right)\right)-\log S+\gamma_{2}.$ (13)
From Lemma 1, we can see that a scaled r.v. $\hat{\lambda}_{\textsc{fpr}}:=\sqrt{S}\lambda_{\textsc{fpr}}$ becomes almost independent of $S$ for large enough $S$. Hence using the change of variables formula, we have that $p_{\lambda_{\textsc{fpr}}}(t)=p_{\hat{\lambda}_{\textsc{fpr}}}(\sqrt{S}t)\sqrt{S}$. Now, evaluating $p_{\lambda_{\textsc{fpr}}}(t)$ at $t=\tau(S)$, we get $p_{\lambda_{\textsc{fpr}}}(t)=p_{\hat{\lambda}_{\textsc{fpr}}}(\gamma_{1})\sqrt{S}$, where the $p_{\hat{\lambda}_{\textsc{fpr}}}(\gamma_{1})$ is independent of $S$. Therefore, ignoring additive constants,
$\log(\textsc{tpr}-\textsc{fpr})$ $\approx\log\sqrt{S}-\log S=\frac{1}{2}\log S-\log S=-\frac{1}{2}\log S.$ (14)
This shows that $\log(\textsc{tpr}-\textsc{fpr})$ depends linearly on $\log S$. ∎
Compared to the model in Section 3.3, our theoretical model predicts $\beta_{S}=-1/2$. Our empirical estimates match this for high fpr. For lower FPR, $\beta_{S}$ decreases, which indicates there may be finer details of $p_{\lambda_{\textsc{fpr}}}$ we are missing in this analysis.
5 MIA Vulnerability For Data Samples
This section extends our investigation from the data-set-level properties (Section 3) to individual data-sample-level properties. Previously Yu et al. (2023) have studied correlations between classes and privacy risk before, but in comparison to their study, we look at data samples independent from classes. We investigate two properties that could potentially influence the vulnerability of data points:
  • Per-sample gradients: Per-sample gradients form the basis for DP individual accounting (Feldman and Zrnic, 2021; Koskela et al., 2023). Under individual Gaussian DP (GDP, Dong et al., 2022) accounting (Koskela et al., 2023), the cumulative $\mu$-GDP privacy loss of individual $i$ with individual gradient norms $g_{i,t}$ is $\mu_{\text{i,tot}}\propto\sqrt{\sum_{t}g_{i,t}^{2}}$. Our hypothesis is that this similarly holds also for MIA vulnerability in non-DP settings.
  • Feature space distances: The pre-trained feature extractors project images into a feature space in which one can compute distances between individual data samples. Our hypothesis is that data samples that are further away from other data samples are easier to distinguish.
5.1 Experimental Setup
Our model training is done as in Section 3.1. The target models are fine-tuned pre-trained ViT-B (Head) and R-50 (Head) models. Below we describe how we measure the sample properties and individual tpr at fpr.
Recording of gradient norms During the training of the target models we record the per-sample gradient norm for every sample at every step (independent of being part of the current batch) efficiently using Opacus (Yousefpour et al., 2021). The recording of the gradient norms is identical to the DP individual accounting, with the exception that we record the gradient norms of out-samples as well. In our analysis, we are utilising the square root of the sum of squared gradient norm per sample over deciles of steps (i.e. the sum of squared individual gradient norms over the first 10% of steps, second 10% of steps, …, last 10% of steps).
$G_{i}^{(\alpha)}=\sqrt{\sum_{t=\lfloor T(\alpha-0.1)\rfloor}^{\lfloor T\alpha\rfloor}g_{i,t}^{2}},\;\alpha=(0.1,0.2,\ldots,1.0)$ (15)
Computation of distance in feature space All samples are projected into a feature space using the pre-trained model. We measure the Euclidean distance of all examples (in and out) to all in-samples in the feature space and report the distance to the nearest in-sample.
tpr at fpr for individual samples We study individual samples, but aggregate the results over multiple data samples by binning the samples based on their data sample properties to reduce the noise. For each data sample we compute its property and assign the sample to a bin based on its percentile in relation to other samples. We then compute the tpr at fpr for each bin. We finally report for each data property bin the percentile of the tpr at fpr.
Throughout our experiments we noticed that the amount of data points needs to be sufficiently large, in order to allow for an appropriate estimation of the tpr at fpr. Thus, we use a sufficient amount of shots ($S$) so that the size of in- and out-set combined is at least $10002|\mathcal{D}|=$ 1000 (see Table A10).
5.2 Individual Gradient Norms Are Predictive of Individual MIA Vulnerability
Figure 5 shows the observed relationships between the MIA vulnerability and the average individual gradient norms at $\textsc{fpr}=0.001$. When plotting the data for all deciles of steps at once, we find that the average gradient norms are correlated with the vulnerability in all our experiments, but the strength of the correlation depends on the data set. The $R^{2}$ (coefficient of determination) compares the observed relationship to a scenario where individual gradient norms would be a perfect predictor (corresponding to a scenario where all points are on the diagonal line). See more results in Table A10 and Figure A.6 in the Appendix.
Figure 5: MIA vulnerability (tpr at fpr; percentiles among a seed) as a function of individual gradient norms (binned among a seed). Data obtained over six seeds. The $R^{2}$ score compares the observed relationship to a scenario where individual gradient norms would be a perfect predictor.
Which parts of the training are more predictive?
In Figure 6, we examine whether the gradient norms at different stages of training are a better predictor of vulnerability by computing the ratio between the $R^{2}$ per iteration decile and $R^{2}$ over the whole training. For most data sets the ratio is constantly at a value of roughly 1, which suggests that gradient norms over even short segments can be as predictive as the norms of the whole trace.
Furthermore, there are deviations that indicate that for some data sets, gradient norms during the early iterations can be relatively more predictive than the norms of the entire trace. This is interesting, as the DP bound would directly depend on the norms of the entire trace.
Figure 6: Ratio between $R^{2}$ per iteration decile and $R^{2}$ over the whole training. A ratio $>1$ means that the decile is as more predictive than whole trace and a ratio $<1$ means that it is less predicitive.
5.3 Distance to Neighbors in Feature Space of Pre-trained Model
Figure 7 shows the observed relationships between the MIA vulnerability and distance to the nearest neighbor in the feature space. The distance of data samples to their nearest neighbor correlates with the vulnerability all analysed data sets but Patch Camelyon and Resics45. The observation holds over all fpr and for R-50 (Head) as well, refer to Figures A.9 and A.10 in the Appendix.
Figure 7: MIA vulnerability (tpr at fpr; percentiles among a seed) as a function of feature space distance (binned among a seed). The results have been obtained over six seeds.
6 Discussion
Under the GDPR and similar legal regimes, machine learning (ML) models trained on personal data that memorise the data are personal data and need to be carefully protected. Our work analyses in which cases trained models would most likely be personal data and in which cases they might not be. This will help in evaluating the risk of different kinds of models, favouring less risky models when possible and paying extra attention to more risky cases.
As the best means of protecting privacy, DP, reduces the utility of models, it is important to understand when it is necessary. Our results highlight that models are most vulnerable to MIA when the number of examples per class is low. This leads to an unfortunate dilemma: DP seems most important in the regime where it likely reduces the model utility the most.
Our model singles out the number of examples per class $S$ as the key predictor of model vulnerability. While our results are based on balanced data sets, it seems likely that minority classes and outliers would be more vulnerable as predicted by our model given their smaller $S$. Studying the vulnerability in unbalanced data and in the presence of outliers is an important area of future work.
Limitations
Our work measures the vulnerability through running one particular MIA. There is no guarantee that models that are not vulnerable to this attack might not be vulnerable to some stronger attack. Formal bounds on attack susceptibility would require something like DP.
Due to computational constraints, our predictive model is based only on transfer learning with fine-tuning only a part of the model. Limited comparison of our model predictions to previous results indicates it may underestimate their vulnerability.
Broader Impact
Our work systematically studies factors influencing privacy risk of trained ML models. This has significant positive impact on users training ML models on personal data by allowing them to understand and limit the risks.
Reproducibility
We provide the code for reproducing the experiments in the supplementary material and will publish the code in an open repository after acceptance of the paper. All used pre-trained models and data sets are publicly available.
Acknowledgements
This work was supported by the Research Council of Finland (Flagship programme: Finnish Center for Artificial Intelligence, FCAI, Grant 356499 and Grant 359111), the Strategic Research Council at the Research Council of Finland (Grant 358247) as well as the European Union (Project 101070617). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the granting authority can be held responsible for them. The authors wish to thank the Finnish Computing Competence Infrastructure (FCCI) for supporting this project with computational and data storage resources. We thank John F. Bronskill for helpful discussions regarding few-shot learning.
Appendix A Training details
A.1 Parameterization
We utilise pre-trained feature extractors BiT-M-R50x1 (R-50) (Kolesnikov et al., 2020) with 23.5M parameters and Vision Transformer ViT-Base-16 (ViT-B) (Dosovitskiy et al., 2021) with 85.8M parameters, both pretrained on the ImageNet-21K data set (Russakovsky et al., 2015). We download the feature extractor checkpoints from the respective repositories.
Following Tobaben et al. (2023) that show the favorable trade-off of parameter-efficient fine-tuning between computational cost, utility and privacy even for small data sets, we only consider fine-tuning subsets of all feature extractor parameters. We consider the following configurations:
  • Head: We train a linear layer on top of the pre-trained feature extractor.
  • FiLM: In addition to the linear layer from Head, we fine-tune parameter-efficient FiLM (Perez et al., 2018) adapters scattered throughout the network. While a diverse set of adapters has been proposed, we utilise FiLM as it has been shown to be competitive in prior work (Shysheya et al., 2023; Tobaben et al., 2023).
A.1.1 Licenses and Access
The licenses and means to access the model checkpoints can be found below.
A.2 Hyperparameter tuning
[…]
Table A1: Hyperparameter ranges used for the Bayesian optimization with Optuna.
lower bound upper bound
batch size 10 $|\mathcal{D}|$
clipping norm 0.2 10
epochs 1 200
learning rate 1e-7 1e-2
A.3 Data sets
Table A2 shows the data sets used in the paper. We base our experiments on a subset of the the few-shot benchmark VTAB (Zhai et al., 2019) that achieves a classification accuracy $>80\%$ and thus would considered to be used by a practitioner. Additionally, we add CIFAR10 which is not part of the original VTAB benchmark.
Table A2: Used datasets in the paper, their minimum and maximum shots $S$ and maximum number of classes $C$ and their test accuracy when fine-tuning a non-DP ViT-B Head. The test accuracy for EuroSAT and Resics45 is computed on the part of the training split that is not used for training the particular model due to both datasets missing an official test split. Note that LiRA requires $2S$ for training the shadow models and thus $S$ is smaller than when only performing fine-tuning.
dataset (max.) min. max. test accuracy test accuracy
$C$ $S$ $S$ (min. $S$) (max. $S$)
Patch Camelyon (Veeling et al., 2018) 2 256 65536 82.8% 85.6%
CIFAR10 (Krizhevsky, 2009) 10 8 2048 92.7% 97.7%
EuroSAT (Helber et al., 2019) 10 8 512 80.2% 96.7%
Pets (Parkhi et al., 2012) 37 8 32 82.3% 90.7%
Resics45 (Cheng et al., 2017) 45 32 256 83.5% 91.6%
CIFAR100 (Krizhevsky, 2009) 100 16 128 82.2% 87.6%
A.3.1 Licenses and Access
The licenses and means to access the data sets can be found below. We downloaded all data sets from TensorFlow datasets https://www.tensorflow.org/datasets but Resics45 which required manual download.
A.4 Compute Resources
All experiments but the R-50 (FiLM) experiments are run on CPU with 8 cores and 16 GB of host memory. The training time depends on the model (ViT is cheaper than R-50), number of shots $S$ and the number of classes $C$ but ranges for the training of one model from some minutes to an hour. This assumes that the images are passed once through the pre-trained backbone and then cached as feature vectors. The provided code implements this optimization.
In Section 5 we record the per-sample gradients during the training. This makes the training much more expensive despite the use of efficient per-sample gradient computation using Opacus Yousefpour et al. (2021). The training of the considered models does still not exceed one hour per model as we only consider settings with relatively low $S$ for this part of the experiments.
The R-50 (FiLM) experiments are significantly more expensive and utilise a NVIDIA V100 with 40 GB VRAM, 10 CPU cores and 64 GB of host memory. The training of 257 shadow models then does not exceed 24h for the settings that we consider.
We estimate that in total we spend around 7 days of V100 and some dozens of weeks of CPU core time but more exact measurements are hard to make.
Appendix B Additional results
In this section, we provide tabular results for our experiments and additional figures that did not fit into the main paper.
B.1 Additional results for Section 3
This Section contains additional results for Section 3.
B.1.1 Vulnerability as a function of shots
This section displays additional results to Figure 1(a) for fpr $\in\{0.1,0.01,0.001\}$ for ViT-B and R-50 in in Figures A.1, A3 and A4.
(a) ViT-B Head tpr at $\textsc{fpr}=0.1$
(b) R-50 Head tpr at $\textsc{fpr}=0.1$
(c) ViT-B Head tpr at $\textsc{fpr}=0.01$
(d) R-50 Head tpr at $\textsc{fpr}=0.01$
(e) ViT-B Head tpr at $\textsc{fpr}=0.001$
(f) R-50 Head tpr at $\textsc{fpr}=0.001$
Figure A.1: MIA vulnerability as a function of shots (examples per class) when attacking a pre-trained ViT-B and R-50 Head trained without DP on different downstream data sets. The errorbars display the minimum and maximum Clopper-Pearson CIs over six seeds and the solid line the median.
Table A3: Median MIA vulnerability over six seeds as a function of $S$ (shots) when attacking a Head trained without DP on-top of a ViT-B. The ViT-B is pre-trained on ImageNet-21k.
dataset classes ($C$) shots ($S$) tpr@fpr=0.1 tpr@fpr=0.01 tpr@fpr=0.001 Patch Camelyon (Veeling et al., 2018) 2 256 0.266 0.086 0.032 512 0.223 0.059 0.018 1024 0.191 0.050 0.015 2048 0.164 0.037 0.009 4096 0.144 0.028 0.007 8192 0.128 0.021 0.005 16384 0.118 0.017 0.003 32768 0.109 0.014 0.002 65536 0.105 0.012 0.002 CIFAR10 (Krizhevsky, 2009) 10 8 0.910 0.660 0.460 16 0.717 0.367 0.201 32 0.619 0.306 0.137 64 0.345 0.132 0.067 128 0.322 0.151 0.082 256 0.227 0.096 0.054 512 0.190 0.068 0.032 1024 0.168 0.056 0.025 2048 0.148 0.039 0.013 EuroSAT (Helber et al., 2019) 10 8 0.921 0.609 0.408 16 0.738 0.420 0.234 32 0.475 0.222 0.113 64 0.400 0.159 0.074 128 0.331 0.155 0.084 256 0.259 0.104 0.049 512 0.213 0.080 0.037 Pets (Parkhi et al., 2012) 37 8 0.648 0.343 0.160 16 0.745 0.439 0.259 32 0.599 0.311 0.150 Resics45 (Cheng et al., 2017) 45 32 0.672 0.425 0.267 64 0.531 0.295 0.168 128 0.419 0.212 0.115 256 0.323 0.146 0.072 CIFAR100 (Krizhevsky, 2009) 100 16 0.814 0.508 0.324 32 0.683 0.445 0.290 64 0.538 0.302 0.193 128 0.433 0.208 0.114
Table A4: Median MIA vulnerability over six seeds as a function of $S$ (shots) when attacking a Head trained without DP on-top of a R-50. The R-50 is pre-trained on ImageNet-21k.
dataset classes ($C$) shots ($S$) tpr@fpr=0.1 tpr@fpr=0.01 tpr@fpr=0.001 Patch Camelyon (Veeling et al., 2018) 2 256 0.272 0.076 0.022 512 0.195 0.045 0.011 1024 0.201 0.048 0.011 2048 0.178 0.041 0.010 4096 0.163 0.033 0.008 8192 0.143 0.026 0.006 16384 0.124 0.019 0.004 32768 0.118 0.016 0.003 65536 0.106 0.012 0.002 CIFAR10 (Krizhevsky, 2009) 10 8 0.911 0.574 0.324 16 0.844 0.526 0.312 32 0.617 0.334 0.183 64 0.444 0.208 0.106 128 0.334 0.159 0.084 256 0.313 0.154 0.086 512 0.251 0.103 0.051 1024 0.214 0.082 0.038 EuroSAT (Helber et al., 2019) 10 8 0.846 0.517 0.275 16 0.699 0.408 0.250 32 0.490 0.236 0.121 64 0.410 0.198 0.105 128 0.332 0.151 0.075 256 0.269 0.111 0.056 512 0.208 0.077 0.036 Pets (Parkhi et al., 2012) 37 8 0.937 0.631 0.366 16 0.745 0.427 0.227 32 0.588 0.321 0.173 Resics45 (Cheng et al., 2017) 45 32 0.671 0.405 0.235 64 0.534 0.289 0.155 128 0.445 0.231 0.121 256 0.367 0.177 0.088 CIFAR100 (Krizhevsky, 2009) 100 16 0.897 0.638 0.429 32 0.763 0.549 0.384 64 0.634 0.414 0.269
B.1.2 Vulnerability as a function of the number of classes
This section displays additional results to Figure 1(b) for fpr $\in\{0.1,0.01,0.001\}$ for ViT-B and R-50 in in Figures A.2, A5 and A6.
(a) ViT-B Head tpr at $\textsc{fpr}=0.1$
(b) R-50 Head tpr at $\textsc{fpr}=0.1$
(c) ViT-B Head tpr at $\textsc{fpr}=0.01$
(d) R-50 Head tpr at $\textsc{fpr}=0.01$
(e) ViT-B Head tpr at $\textsc{fpr}=0.001$
(f) R-50 Head tpr at $\textsc{fpr}=0.001$
Figure A.2: MIA vulnerability as a function of $C$ (classes) when attacking a ViT-B and R-50 Head fine-tuned without DP on different data sets where the classes are randomly sub-sampled and $S=32$. The solid line displays the median and the errorbars the min and max clopper-pearson CIs over 12 seeds.
Table A5: Median MIA vulnerability over 12 seeds as a function of $C$ (classes) when attacking a Head trained without DP on-top of a ViT-B. The Vit-B is pre-trained on ImageNet-21k.
dataset shots ($S$) classes ($C$) tpr@fpr=0.1 tpr@fpr=0.01 tpr@fpr=0.001 Patch Camelyon (Veeling et al., 2018) 32 2 0.467 0.192 0.080 CIFAR10 (Krizhevsky, 2009) 32 2 0.494 0.167 0.071 4 0.527 0.217 0.115 8 0.574 0.262 0.123 EuroSAT (Helber et al., 2019) 32 2 0.306 0.100 0.039 4 0.298 0.111 0.047 8 0.468 0.211 0.103 Pets (Parkhi et al., 2012) 32 2 0.232 0.045 0.007 4 0.324 0.092 0.035 8 0.296 0.094 0.035 16 0.406 0.158 0.069 32 0.553 0.269 0.136 Resics45 (Cheng et al., 2017) 32 2 0.272 0.084 0.043 4 0.322 0.119 0.056 8 0.496 0.253 0.148 16 0.456 0.204 0.108 32 0.580 0.332 0.195 CIFAR100 (Krizhevsky, 2009) 32 2 0.334 0.088 0.035 4 0.445 0.150 0.061 8 0.491 0.223 0.121 16 0.525 0.256 0.118 32 0.553 0.276 0.153 64 0.612 0.350 0.211
Table A6: Median MIA vulnerability over 12 seeds as a function of $C$ (classes) when attacking a Head trained without DP on-top of a R-50. The R-50 is pre-trained on ImageNet-21k.
dataset shots ($S$) classes ($C$) tpr@fpr=0.1 tpr@fpr=0.01 tpr@fpr=0.001 Patch Camelyon (Veeling et al., 2018) 32 2 0.452 0.151 0.041 CIFAR10 (Krizhevsky, 2009) 32 2 0.404 0.146 0.060 4 0.560 0.266 0.123 8 0.591 0.318 0.187 EuroSAT (Helber et al., 2019) 32 2 0.309 0.111 0.050 4 0.356 0.144 0.064 8 0.480 0.233 0.123 Pets (Parkhi et al., 2012) 32 2 0.249 0.068 0.029 4 0.326 0.115 0.056 8 0.419 0.173 0.075 16 0.493 0.245 0.127 32 0.559 0.294 0.166 Resics45 (Cheng et al., 2017) 32 2 0.310 0.103 0.059 4 0.415 0.170 0.083 8 0.510 0.236 0.119 16 0.585 0.311 0.174 32 0.644 0.382 0.218 CIFAR100 (Krizhevsky, 2009) 32 2 0.356 0.132 0.054 4 0.423 0.176 0.087 8 0.545 0.288 0.163 16 0.580 0.338 0.196 32 0.648 0.402 0.244 64 0.711 0.476 0.320
B.1.3 Data for FiLM and from scratch training
Table A7: MIA vulnerability data used in Figure 3(b). Note that the data from Carlini et al. (2022) is only partially tabular, thus we estimated the tpr at fpr from the plots in the Appendix of their paper.
model dataset classes shots source tpr@ tpr@ tpr@ ($C$) ($S$) fpr=0.1 fpr=0.01 fpr=0.001 R-50 FiLM CIFAR10 10 50 This work 0.482 0.275 0.165 (Krizhevsky, 2009) CIFAR100 100 10 Tobaben et al. (2023) 0.933 0.788 0.525 (Krizhevsky, 2009) 25 Tobaben et al. (2023) 0.766 0.576 0.449 50 Tobaben et al. (2023) 0.586 0.388 0.227 100 Tobaben et al. (2023) 0.448 0.202 0.077 EuroSAT 10 8 This work 0.791 0.388 0.144 (Helber et al., 2019) Patch Camelyon 2 256 This work 0.379 0.164 0.076 (Veeling et al., 2018) Pets 37 8 This work 0.956 0.665 0.378 (Parkhi et al., 2012) Resics45 45 32 This work 0.632 0.379 0.217 (Cheng et al., 2017) from scratch CIFAR10 10 2500 Carlini et al. (2022) 0.300 0.110 0.084 (Krizhevsky, 2009) (wide ResNet) CIFAR100 100 250 Carlini et al. (2022) 0.700 0.400 0.276 (Krizhevsky, 2009)
B.1.4 Predicting data set vulnerability as function of $S$ and $C$
This section provides additional results for the model based on Equation 2
Table A8: Results for fitting Equation A1 with statsmodels Seabold and Perktold (2010) to ViT Head data at fpr $\in\{0.1,0.01,0.001,0.0001,0.00001\}$. We utilize an ordinary least squares. The test $R^{2}$ assesses the fit to the data of R-50 Head.
coeff. fpr $R^{2}$ test $R^{2}$ coeff. value std. error $t$ $p>|z|$ coeff. $[0.025$ coeff. $0.975]$ $\beta_{S}$ (for $S$) 0.1 0.952 0.907 -0.506 0.011 -44.936 $0$0.529 -0.484 0.01 0.946 0.854 -0.555 0.014 -39.788 $0$0.582 -0.527 0.001 0.930 0.790 -0.627 0.019 -32.722 $0$0.664 -0.589 0.0001 0.852 0.618 -0.741 0.035 -21.467 $0$0.809 -0.673 0.00001 0.837 0.404 -0.836 0.045 -18.690 $0$0.924 -0.748 $\beta_{C}$ (for $C$) 0.1 0.952 0.907 0.090 0.021 4.231 0.000 0.048 0.131 0.01 0.946 0.854 0.182 0.026 6.960 0.000 0.131 0.234 0.001 0.930 0.790 0.300 0.036 8.335 0.000 0.229 0.371 0.0001 0.852 0.618 0.363 0.065 5.616 0.000 0.236 0.491 0.00001 0.837 0.404 0.569 0.085 6.655 0.000 0.400 0.737 $\beta_{0}$ (intercept) 0.1 0.952 0.907 0.314 0.045 6.953 0.000 0.225 0.402 0.01 0.946 0.854 0.083 0.056 1.491 0.137 -0.027 0.193 0.001 0.930 0.790 -0.173 0.077 -2.261 0.025 -0.324 -0.022 0.0001 0.852 0.618 -0.303 0.138 -2.202 0.029 -0.575 -0.032 0.00001 0.837 0.404 -0.615 0.180 -3.414 0.001 -0.970 -0.260
Figure A.3 shows the performance for all considered fpr.
[…]
Figure A.3: Predicted MIA vulnerability as a function of $S$ (shots) using a model based on Equation 2 fitted Table A3 (ViT-B). The triangles show the median tpr for the train set (ViT-B; Table A3) and circle the test set (R-50; Table A4) over six seeds. Note that the triangles and dots for $C=10$ are for EuroSAT.
B.2 Simpler variant of the prediction model
[…]
The general form this variant can be found in Equation A1, where $\beta_{S},\beta_{C}$ and $\beta_{0}$ are the learnable regression parameters.
$\log_{10}(\textsc{tpr})=\beta_{S}\log_{10}(S)+\beta_{C}\log_{10}(C)+\beta_{0}$ (A1)
Table A9 provides tabular results on the performance of the variant.
Table A9: Results for fitting Equation A1 with statsmodels Seabold and Perktold (2010) to ViT Head data at fpr $\in\{0.1,0.01,0.001,0.0001,0.00001\}$. We utilize an ordinary least squares. The test $R^{2}$ assesses the fit to the data of R-50 Head.
coeff. fpr $R^{2}$ test $R^{2}$ coeff. value std. error $t$ $p>|z|$ coeff. $[0.025$ coeff. $0.975]$ $\beta_{S}$ (for $S$) 0.1 0.908 0.764 -0.248 0.008 -30.976 $0$0.264 -0.233 0.01 0.940 0.761 -0.416 0.011 -36.706 $0$0.438 -0.393 0.001 0.931 0.782 -0.553 0.017 -32.507 $0$0.586 -0.519 0.0001 0.865 0.628 -0.697 0.031 -22.274 $0$0.758 -0.635 0.00001 0.862 0.400 -0.802 0.040 -20.311 $0$0.880 -0.725 $\beta_{C}$ (for $C$) 0.1 0.908 0.764 0.060 0.015 3.955 0.000 0.030 0.089 0.01 0.940 0.761 0.169 0.021 7.941 0.000 0.127 0.211 0.001 0.931 0.782 0.297 0.032 9.303 0.000 0.234 0.360 0.0001 0.865 0.628 0.371 0.059 6.328 0.000 0.255 0.486 0.00001 0.862 0.400 0.580 0.076 7.679 0.000 0.431 0.729 $\beta_{0}$ (intercept) 0.1 0.908 0.764 0.029 0.032 0.913 0.362 -0.034 0.093 0.01 0.940 0.761 -0.118 0.045 -2.613 0.010 -0.208 -0.029 0.001 0.931 0.782 -0.295 0.068 -4.345 $0$0.429 -0.161 0.0001 0.865 0.628 -0.387 0.125 -3.104 0.002 -0.633 -0.141 0.00001 0.862 0.400 -0.683 0.159 -4.288 $0$0.996 -0.369
Figure A.4 plots the performance of the variant similar to Figure 3(a) in the main text.
(a) tpr at $\textsc{fpr}=0.1$
(b) tpr at $\textsc{fpr}=0.01$
(c) tpr at $\textsc{fpr}=0.001$
(d) tpr at $\textsc{fpr}=0.0001$
(e) tpr at $\textsc{fpr}=0.00001$
Figure A.4: Predicted MIA vulnerability as a function of $S$ (shots) using a model based on Equation A1 fitted Table A3 (ViT-B). The triangles show the median tpr for the train set (ViT-B; Table A3) and circle the test set (R-50; Table A4) over six seeds. Note that the triangles and dots for $C=10$ are for EuroSAT.
B.3 Additional results for Section 4
This section provides additional results for Section 4.
B.4 Rates of In- and Out-distribution Parameters as Function of data set Properties
LiRA assumes the attacker can build Gaussian distributions of losses for each data point $x$ through the training of shadow models (see Section 2). We denote the distributions as following:
  • In-distribution: Distribution of the losses of shadow models using $x$ for training ($x\in\mathcal{D}$)
  • Out-distribution: Distribution of the losses of shadow models not using $x$ for training ($x\notin\mathcal{D}$)
In Figure A.5 we visualise the changes of the two distributions as functions of $S$ and $C$.
(a) Distribution means
(b) Distribution variances
(c) Distribution standard deviations
(d) Difference of Means
(e) Difference of Means / std
(f) Difference of Means / var
Figure A.5: Means of observed statistics of the in- and out-distributions for the simplified model with in-class $\sigma=0.1$. The statistics are obtained for $C=2$.
B.5 Additional results for Section 5
This section provides additional results for Section 5. Figures A.6 and A.7 provide additional data for all considered fpr and for both ViT-B (Head) and R-50 (Head). Table A10 the $R^{2}$ scores for ViT-B (Head) and the experiment settings for ViT-B (Head) and R-50 (Head). Figure A.8 the relative $R^{2}$ scores for R-50 (Head).
B.5.1 MIA vulnerability as a function of individual gradient norms
Figure A.6: ViT-B (Head): MIA vulnerability (tpr at fpr; percentiles among a seed) as a function of individual gradient norms (binned among a seed). The results have been obtained over six seeds. The number in the bottom right is the $R^{2}$ (coefficient of determination) score when comparing the observed relationship between tpr at fpr and individual gradient norms to a scenario where individual gradient norms are a perfect predictor of individual MIA vulnerability (corresponding to a scenario where all points are on the diagonal line).
Figure A.7: R-50 (Head): MIA vulnerability (tpr at fpr; percentiles among a seed) as a function of individual gradient norms (binned among a seed). The results have been obtained over six seeds. The number in the bottom right is the $R^{2}$ (coefficient of determination) score when comparing the observed relationship between tpr at fpr and individual gradient norms to a scenario where individual gradient norms are a perfect predictor of individual MIA vulnerability (corresponding to a scenario where all points are on the diagonal line).
Table A10: ViT-B (Head): $R^{2}$ (coefficient of determination) score when comparing the observed relationship between tpr at fpr and individual gradient norms to a scenario where individual gradient norms are a perfect predictor of individual MIA vulnerability (corresponding to a scenario where all points are on the diagonal line).
dataset shots ($S$) classes ($C$) data set size (|$\mathcal{D}$|) $R^{2}$ at fpr=0.1 $R^{2}$ at fpr=0.01 $R^{2}$ at fpr=0.001 Patch Camelyon (Veeling et al., 2018) 256 2 512 0.98 0.97 0.81 CIFAR10 (Krizhevsky, 2009) 64 10 640 0.98 0.98 0.92 EuroSAT (Helber et al., 2019) 64 10 640 0.98 0.97 0.92 Pets (Parkhi et al., 2012) 32 37 1184 0.95 0.93 0.89 Resics45 (Cheng et al., 2017) 32 45 1440 0.95 0.97 0.97 CIFAR100 (Krizhevsky, 2009) 16 100 1600 0.89 0.89 0.86
Figure A.8: R-50 (Head): Ratio between $R^{2}$ per iteration decile and $R^{2}$ over the whole training. A ratio $>1$ means that the decile is as more predictive than whole trace and a ratio $<1$ means that it is less predictive.
B.5.2 MIA vulnerability as a function of feature space distance
This section provides additional data regarding the individual MIA prediction using feature space distances.
Figure A.9: ViT-B (Head): MIA vulnerability (tpr at fpr; percentiles among a seed) as a function of feature space distance (binned among a seed). The results have been obtained over six seeds.
Figure A.10: R-50 (Head): MIA vulnerability (tpr at fpr; percentiles among a seed) as a function of feature space distance (binned among a seed). The results have been obtained over six seeds.
Appendix C Missing proofs of Section 4.2
See 1
Proof.
Recall
fpr $=\Pr\left(\underbrace{-\frac{1}{2\sigma^{2}}(\mu_{\textrm{in}}-\mu_{\textrm{out}})\left[-2t_{x}+(\mu_{\textrm{in}}+\mu_{\textrm{out}})\right]}_{\lambda_{\textsc{fpr}}}>\tau\mid x\notin\mathcal{D}_{\textrm{target}}\right)$ (A2)
tpr $=\Pr\left(\underbrace{-\frac{1}{2\sigma^{2}}(\mu_{\textrm{in}}-\mu_{\textrm{out}})\left[-2t_{x}+(\mu_{\textrm{in}}+\mu_{\textrm{out}})\right]}_{\lambda_{\textsc{tpr}}}>\tau\mid x\in\mathcal{D}_{\textrm{target}}\right).$ (A3)
We denote the class means for the $m$th in/out shadow data sets with $r_{y,m}^{(\textrm{in})}$ and $r_{y,m}^{(\textrm{out})}$
$r_{y,m}^{(\textrm{in})}$ $=\frac{1}{S}\left(\sum_{i=1}^{S-1}x_{i,m,y}+x\right),\;r_{y,m}^{(\textrm{out})}=\frac{1}{S}\left(\sum_{i=1}^{S}x_{i,m,y}\right),$ (A4)
where $x_{i,m,y}$ denotes the $i$th element of the $m$th shadow data set with label $y$. The corresponding scores $s_{y,m}^{(\textrm{in})}=\langle x,r_{y,m}^{(\textrm{in})}\rangle$ and $s_{y,m}^{(\textrm{out})}=\langle x,r_{y,m}^{(\textrm{out})}\rangle$ which can be written as
$s_{y,m}^{(\textrm{in})}$ $=\frac{1}{S}\left(\langle x,\sum_{i=1}^{S-1}x_{i,m,y}\rangle+\langle x,x\rangle\right)=\frac{1}{S}\left(\langle x,\sum_{i=1}^{S}x_{i,m,y}\rangle+\langle x,x-x_{S,m,y}\rangle\right)$ (A5)
$s_{y,m}^{(\textrm{out})}$ $=\frac{1}{S}\left(\langle x,\sum_{i=1}^{S}x_{i,m,y}\rangle\right)$ (A6)
Now, assuming we have arbitrarily many shadow models, we can write the LiRA statistics as
$\mu_{\textrm{in}}=\operatorname{\mathbb{E}}_{D_{m}}[s_{y,m}^{(\textrm{in})}],\;\mu_{\textrm{out}}=\operatorname{\mathbb{E}}_{D_{m}}[s_{y,m}^{(\textrm{out})}]=\langle x,m_{y}\rangle$ (A7)
$\Rightarrow$ $\mu_{\textrm{in}}-\mu_{\textrm{out}}=\frac{1}{S}\langle x,x-\mathbb{E}_{D_{m}}[x_{S,m}]\rangle=\frac{1}{S}\langle x,x-m_{y}\rangle,$ (A8)
where $m_{y}$ is the true center of the $y$th cluster. For $\sigma$ we have
$\sigma^{2}$ $=\operatorname{Var}(s_{y,m}^{(\textrm{out})})$ (A9)
$=\frac{1}{S^{2}}\sum_{i=1}^{S}\operatorname{Var}(\langle x,x_{i,m}\rangle)$ (A10)
$=\frac{1}{S}\operatorname{Var}(\langle x,x_{m}^{(j)}\rangle)$ (A11)
$=\frac{1}{S}\operatorname{Var}(x_{m}^{(j)})||x||^{2}=\frac{1}{S}s^{2}||x||^{2},$ (A12)
where $x_{m}^{(j)}$ denotes the $j$th dimension of the vector $x_{m}$ and $s$ is the std of each dimension of $x_{m}$. Now we can write the $\lambda_{\textsc{fpr}}$ as
$\lambda_{\textsc{fpr}}=-\frac{1}{2s^{2}||x||^{2}}\langle x,x-m_{y}\rangle\left(-2t_{x}+2\langle x,m_{y}\rangle+\frac{1}{S}\langle x,x-m_{y}\rangle\right).$ (A13)
In the fpr case, we have that $x\notin\mathcal{D}_{\textrm{target}}$, therefore for the $t_{x}$ we have
$t_{x}=\langle x,r_{y}^{\textrm{(target)}}\rangle\sim\langle x,m_{y}\rangle+\frac{s}{\sqrt{S}}\langle x,z\rangle,$ (A14)
where $z\sim N(0,I_{d})$. Substituting this into $\lambda_{\textsc{fpr}}$ we have
$\lambda_{\textsc{fpr}}=-\frac{1}{2s^{2}||x||^{2}}\langle x,x-m_{y}\rangle\left(\frac{2s}{\sqrt{S}}\langle x,z\rangle+\frac{1}{S}\langle x,x-m_{y}\rangle\right).$ (A15)
We see that $\lambda_{\textsc{fpr}}$ evolves as $O(1/\sqrt{S})$.
Next, we will focus on the $\lambda_{\textsc{tpr}}$. The expression remains otherwise the same as for $\lambda_{\textsc{fpr}}$, but since $x\in\mathcal{D}_{\textrm{target}}$ the $t_{x}$ changes to
$t_{x}\sim\langle x,m_{y}\rangle+\frac{1}{S}\langle x,x\rangle+s\frac{1}{\sqrt{S}}\sqrt{1-\frac{1}{S}}\langle x,z\rangle.$ (A16)
Hence, for large enough $S$ we have
$\lambda_{\textsc{tpr}}\approx-\frac{1}{2s^{2}||x||^{2}}\langle x,x-m_{y}\rangle\left(\frac{2s}{\sqrt{S}}\langle x,z\rangle-\frac{2}{S}||x||^{2}+\frac{1}{S}\langle x,x-m_{y}\rangle\right).$ (A17)